# Strangeness production in $O+O$ collisions at $\sqrt{s_{\mathrm{NN}}}=7$ TeV using various model approaches

### v1

@@ 1 Introduction
The predictions presented in this work are based on recently updated EPOS framework (EPOS4) [15, 16, 17] and two different versions of AMPT model [18, 19, 20]. Both models employ different mechanisms for strangeness production. EPOS4 includes the QGP phase in its hydrodynamic evolution with the lattice QCD equation of state(EoS). AMPT, on the other hand, does not have a fully chemically equilibrated system as compared to EPOS4. In this article, we report on the predictions of strange hadron ($\mathrm{K}^{0}_{\mathrm{S}}$, $\Lambda+\overline{\Lambda}$, $\Xi^{-}+\overline{\Xi}^{+}$, $\Omega^{-}+\overline{\Omega}^{+}$, and $\phi$) production in $O+O$ collisions at $\sqrt{s_{\mathrm{NN}}}=7$ TeV using EPOS4 and AMPT. The observables under study include $p_{T}$ spectra, rapidity density distributions ($dN/dy$), and multiplicity dependence of the yield ratios of strange hadrons relative to pions.
@@ 3 Results and Discussions
Figure 2: (Color online) $p_{T}$-spectra of (multi)strange hadrons ($\mathrm{K}^{0}_{\mathrm{S}}$, $\Lambda+\overline{\Lambda}$, $\Xi^{-}+\overline{\Xi}^{+}$, $\Omega^{-}+\overline{\Omega}^{+}$, and $\phi$) from EPOS4 and AMPT in $O+O$ collisions at $\sqrt{s_{\mathrm{NN}}}=7$ TeV for 0–5% centrality classes. $p_{T}$ spectra for different hadrons are scaled by a factor of $10$ for better visualization. In figure, solid lines represent EPOS4, while dashed lines correspond to AMPT-SM, and dotted lines represent AMPT-Def.
@@
Figure 3: (Color online) Multiplicity dependence of the integrated yield ($dN/dy$) at mid-rapidity ($\lvert y\rvert<0.5$) for strange hadrons ($\mathrm{K}^{0}_{\mathrm{S}}$, $\Lambda$, $\Xi$, $\phi$, and $\Omega$) in $O+O$ collisions at $\sqrt{s_{\mathrm{NN}}}=7$ TeV using AMPT-Def, AMPT-SM and EPOS4.

### v2

@@ 1 Introduction
The predictions presented in this work are based on recently updated EPOS framework (EPOS4) [15, 16, 17] and two different versions of AMPT model [18, 19, 20]. Both models employ different mechanisms for strangeness production. EPOS4 includes the QGP phase in its hydrodynamic evolution with the lattice QCD equation of state (EoS). AMPT, on the other hand, does not have a fully chemically equilibrated system as compared to EPOS4. In this article, we report on the predictions of (multi-)strange hadron ($\mathrm{K}^{0}_{\mathrm{S}}$, $\Lambda+\overline{\Lambda}$, $\Xi^{-}+\overline{\Xi}^{+}$, $\Omega^{-}+\overline{\Omega}^{+}$, and $\phi$) production in $O+O$ collisions at $\sqrt{s_{\mathrm{NN}}}=7$ TeV using EPOS4 and AMPT. The observables under study include $p_{T}$ spectra, rapidity density distributions ($dN/dy$), and multiplicity dependence of the yield ratios of strange hadrons relative to pions.
@@ 3 Results and Discussions
Figure 2: (Color online) $p_{T}$-spectra of (multi-)strange hadrons ($\mathrm{K}^{0}_{\mathrm{S}}$, $\Lambda+\overline{\Lambda}$, $\Xi^{-}+\overline{\Xi}^{+}$, $\Omega^{-}+\overline{\Omega}^{+}$, and $\phi$) from EPOS4 and AMPT in $O+O$ collisions at $\sqrt{s_{\mathrm{NN}}}=7$ TeV for 0–5% centrality classes. $p_{T}$ spectra for different hadrons are scaled by a factor of 1010 for better visualization. In figure, solid lines represent EPOS4, while dashed lines correspond to AMPT-SM, and dotted lines represent AMPT-Def.
@@
Figure 3: (Color online) Multiplicity dependence of the integrated yield ($dN/dy$) at mid-rapidity $\mid y\mid<0.5$ for (multi-)strange hadrons ($\mathrm{K}^{0}_{\mathrm{S}}$, $\Lambda$, $\Xi$, $\phi$, and $\Omega$) in $O+O$ collisions at $\sqrt{s_{\mathrm{NN}}}=7$ TeV using AMPT-Def, AMPT-SM and EPOS4.